# Online Transfer Learning for RSV Case Detection

### 03 Feb 2024 v1

@@ Abstract
Transfer learning has become a pivotal technique in machine learning, renowned for its effectiveness in various real-world applications. However, a significant challenge arises when applying this approach to sequential epidemiological data, often characterized by a scarcity of labeled information. To address this challenge, we introduce Predictive Volume-Adaptive Weighting (PVAW), a novel online multi-source transfer learning method. PVAW innovatively implements a dynamic weighting mechanism within an ensemble model, allowing for the automatic adjustment of weights based on the relevance and contribution of each source and target model. We demonstrate the effectiveness of PVAW through its application in analyzing Respiratory Syncytial Virus (RSV) data, collected over multiple seasons at the University of Pittsburgh Medical Center. Our method showcases significant improvements in model performance over existing baselines, highlighting the potential of online transfer learning in handling complex, sequential data. This study not only underscores the adaptability and sophistication of transfer learning in healthcare but also sets a new direction for future research in creating advanced predictive models.
@@
online transfer learning, ensemble method, dynamic weighting mechanism, electronic health record, RSV case detection
@@ I Introduction
Machine learning has made substantial advancements in recent decades, with its applications spanning a wide range of fields such as image and speech recognition, natural language processing, and autonomous driving. Despite these achievements, machine learning in biomedicine faces significant challenges, particularly in data collection. The acquisition of labeled data can be very costly or even unfeasible due to factors like ethical considerations, patient privacy, and the scarcity of certain diseases. These challenges have led researchers to increasingly rely on utilizing data from related domains that have a more abundant supply of data.
@@
To better leverage knowledge from these related yet distinct domains, transfer learning has emerged as a promising approach to enhance model performance. Fundamental to transfer learning are the concepts of the ‘source domain’ and ‘target domain’. The source domain, usually rich in data, serves as the foundation for knowledge extraction. It is characterized by its abundance of well-labeled datasets, which aid models in understanding and encoding complex patterns. In contrast, the target domain, where this pre-learned knowledge is applied, often features a more limited dataset. Transfer learning bridges this gap by adapting models trained on the source domain to function effectively in the target domain, thereby helping address the challenges associated with data scarcity.
@@
Traditional transfer learning approaches [1, 2, 3] typically depend on having complete datasets from both the source and target domains at the start of the training process. However, this is often not feasible, especially in dynamic environments where target domain data is collected progressively. To tackle this issue, the concept of online transfer learning [4, 5, 6] has been introduced. This method is designed for scenarios where target domain data is acquired incrementally over time, while source domain data is already accessible before training begins. In such cases, transferring knowledge from the source domain becomes crucial, particularly because the limited initial data in the target domain may be insufficient for effective learning. The extensive and diverse information available from the source domains can significantly compensate for this shortfall, providing a foundational knowledge base that the model can build upon as more target domain data becomes available. Therefore, the efficiency and effectiveness of learning in the target domain are greatly enhanced by the transferred knowledge from the source domains.
@@
Online transfer learning entails leveraging knowledge from a static source domain and applying it to an ongoing, evolving target task. The model is incrementally updated as new data are received. A crucial element of this process is the utilization of a loss function to fine-tune the model in accordance with the latest incoming data, thus ensuring its continued relevance and efficiency. This adaptability renders online transfer learning particularly suitable for real-time applications, effectively handling issues such as concept drift and providing a more versatile approach to knowledge transfer.
@@
In traditional transfer learning, domains are typically categorized based on similar samples, such as distinguishing between images of cats and dogs, which allows for the straightforward grouping of domains with analogous characteristics. However, online transfer learning presents a nuanced challenge as it focuses on the evolution of a single task over time, making the determination of where to apply temporal segmentation less clear. This difficulty arises because the task itself remains constant, but the data representing the task may change significantly due to temporal factors. This complexity is compounded by the inherent challenge of understanding the relationships among temporally segmented domains. Unlike static categorizations, where similarities and differences are often clear and measurable, the dynamic nature of temporal domains means that the relevance of past data to the present task can be fluid and elusive. For instance, in healthcare, temporal segmentation might involve dividing patient data into annual periods to capture seasonal disease trends. With multiple source domains, integrating them optimally into an ensemble model is challenging. A critical question arises: how do we determine which source domains are most relevant to the current period (or target domain) and which are less relevant or unrelated?
@@
Our proposed method employs a sequential data acquisition strategy to tackle this challenge. Consider a healthcare scenario where the target domain is predicting patient outcomes in a new medical facility. To address initial data scarcity in this facility, our approach significantly relies on pre-trained models derived from source domain data, such as historical patient outcomes from established facilities. These pre-trained source models form the foundation of our ensemble model. This ensemble is adept at autonomously evaluating and determining the relevance of each source model’s learned patterns to the current patient data in the new facility. For instance, a source model trained on a large urban hospital might be initially prioritized for its broad spectrum of cases. As more patient data accumulates in the new facility, the ensemble dynamically adjusts its reliance on this urban model, potentially shifting focus to another source model that aligns more closely with emerging local trends. This strategic utilization of pre-trained source models ensures that the most effective ones are prioritized during the critical early stages of training in the target domain.
@@
As the target domain accumulates more data, the ensemble model modifies its approach. Initially dependent on pre-trained source models, it gradually begins to prioritize the continually updated target model. This change is twofold: a shift in reliance from source to target models and a nuanced adjustment in the weight each model holds in decision-making. This transition enables the model to evolve alongside the growing target domain dataset. The distinct roles of shifting focus and adjusting weights demonstrate the adaptability of our online transfer learning approach, particularly in environments where data is collected over time.
@@
Our framework makes two main contributions. First, we introduce an online multi-source transfer learning architecture to study biomedical information over multiple seasons. Second, we explore the performance of different ways of generating weights for the ensemble model to improve its accuracy.
@@
We evaluate the effectiveness of our approach using real-world Respiratory Syncytial Virus (RSV) data that were collected over multiple seasons from patient encounters in Emergency Departments (EDs) at the University of Pittsburgh Medical Center (UPMC). Our primary goal is to utilize reported signs and symptoms to accurately classify patients as either having RSV or not, a critical step towards improving diagnosis and treatment strategies. RSV is a major cause of respiratory infections that lead to hospitalizations [7]. The burden of RSV continues to grow, especially in young children aged less than 5 years old[8, 9]. It has been also identified as an important disease causing hospitalizations of the elderly and high-risk adults [10].
Characteristically, RSV exhibits a seasonal pattern in case numbers [11]. In our approach, we specifically segment the RSV case data into distinct domains based on a defined annual cycle, with each domain spanning from June of the previous year to May of the current year. This segmentation strategy is crucial to our framework, aiming to enhance the predictive accuracy of RSV case detection by incorporating data from multiple preceding seasons as source inputs. For instance, in our experiments, we use patient encounter data ranging from June 2011 to May 2018 to develop an ensemble model capable of classifying encounters in the period from June 2018 to May 2019. The experimental results support that our approach improves predictive performance, particularly related to public health planning and resource allocation for anticipated RSV seasons.
@@ II Related Work
Traditional transfer learning methodologies typically operate under the premise that the source and target domains are predefined, with the primary objective being the transference of knowledge from the source to the target domain. Numerous comprehensive surveys [12, 13, 14, 15, 16, 17] have thoroughly examined and classified a broad range of transfer learning approaches, considering factors like algorithm types and their various use cases. A notable recent development is the growing body of literature reviews focusing on the integration of transfer learning with other advanced machine learning techniques. These include combinations with deep learning [18, 19], reinforcement learning [20, 21, 22], and federated learning [23, 24]. Additionally, some reviews have specifically targeted the application of transfer learning in key areas such as computer vision (CV) [25, 26, 27, 28], natural language processing (NLP) [29, 30, 31], medical image analysis [32, 33], and wireless communication [34, 35], providing insights into the practical implications and utility of transfer learning in these specific fields.
@@
Online transfer learning has emerged to address situations where data are received sequentially and are not available at the start of training. This approach is particularly relevant in dynamic environments where real-time data processing is crucial. Some popular online transfer learning methods have demonstrated advanced performance in many real-world scenarios. Online transfer learning methods can be categorized into data-based methods and model-based methods.
@@
Data-based transfer learning methods involve the transfer of either the original data or their transformed features to a target domain, allowing the target models to be trained with these augmented data, thereby enriching the available data within the target domain. Online Heterogeneous Knowledge Transition (OHKT) [36] initialize the target model using co-occurrence data with pseudo labels generated from the source model.
@@
Conversely, model-based transfer learning methods focus on building models within the source domain, adapting both the model structure and parameters from this domain. Here, ‘model structure’ refers to the architecture or configuration of the model (such as layers in a neural network), while ‘model parameters’ are the specific weights and biases that are learned from the source data. The target models utilize these insights from the source domain, with the aim of potentially enhancing performance. Online Heterogeneous Transfer by Hedge Ensemble (OHTHE) [37] handles the online heterogeneous transfer problem by updating weights of offline and online classifiers based on the Hedge weighting strategy.
@@
For our online transfer learning algorithm, we utilized RSV detection as an illustrative example task. Machine learning algorithms have been increasingly utilized for detecting RSV. Tso et al. [38] developed an extreme gradient boost decision tree model to detect RSV using clinical measurements in electronic health records (EHR), such as respiratory rate and peripheral oxygen saturation, from EDs and inpatient visits in 2019. In a separate study, Cord et al. [39] evaluated various methods, including Logistic Regression, Gaussian Naive Bayes, and Random Forest, before selecting the Light Gradient Boosting Machine algorithm for their RSV detection models.
@@ III-A Problem Formulation
In this study, we address the challenge of binary classification in environments where data are not static but arrive sequentially, often correlated with specific time frames such as seasons. Our focus is on developing a method for studying information collected over seasons using Bayesian networks as the foundation of our classifier training.
Bayesian networks, a type of probabilistic graphical model, are particularly well-suited for this task. They express conditional dependencies between a set of variables through a Directed Acyclic Graph (DAG). In these networks, each node represents a variable, encompassing both input features $\mathbf{x}\in\mathbb{R}^{d}$ and binary labels $y\in\{0,1\}$. The edges denote probabilistic dependencies between these variables. This structure allows Bayesian networks to model complex relationships within the data and to infer the likelihood of certain outcomes given a set of observed variables, thereby making them ideal for classification tasks.
Given a sufficient number of training instances $(\mathbf{x},y)$, Bayesian networks are adept at predicting the probability that a known cause (represented by the feature vector $\mathbf{x}$ was the contributing factor for the observed event (the label $y$). This probabilistic approach not only offers predictions but also provides insights into the strength and nature of the relationships between different features and the outcome.
In real-world applications, each epidemic season’s data can be considered a distinct domain. This temporal characteristic poses a unique challenge, particularly at the start of a new period when data availability is limited. To address this, we propose a transfer learning approach where the current season is treated as the target domain, and data from multiple previous seasons are utilized as source domains. Our method employs data from these multiple source domains, denoted as $D^{S}=\{D_{S}^{1},D_{S}^{2},\ldots,D_{S}^{n}\}$, where each $D_{S}^{i}$ corresponds to a specific previous season’s data, alongside a single target domain $D_{T}$. We assume that patterns and knowledge from previous seasons can be leveraged to effectively predict or analyze trends in the current season, despite the initial scarcity of data. The primary goal is to leverage historical data to inform and improve the predictive model for the current season, under the assumption that underlying patterns and relationships in the data persist over time.
@@ III-B Base Model
Our methodology initiates with the implementation of the Naive Bayes classifier, a cornerstone of probabilistic classification techniques. Central to Naive Bayes is the simplification of probability calculations, achieved by the assumption of conditional independence among features given a class label. This assumption allows the classifier to compute the probability of a class label $y$ given a set of features $x_{1},\ldots,x_{n}$ using the formula:
$P\left(y\mid x_{1},\ldots,x_{n}\right)$ $=\frac{P(y)P\left(x_{1},\ldots,x_{n}\mid y\right)}{P\left(x_{1},\ldots,x_{n}\right)}$ (1)
$=\frac{P(y)\prod_{i=1}^{n}P\left(x_{i}\mid y\right)}{P\left(x_{1},\ldots,x_{n}\right)}.$ (2)
The effectiveness of Naive Bayes lies in its simplicity and its computational and statistical efficiency, especially in cases where the features are reasonably independent given the class label.
@@ III-C Online Multi-source Ensemble Model
In our framework, for each source domain $D_{S}^{i}$, we train a unique offline base model $f_{S}^{i}$ as a classifier on this domain. Similarly, a specific online training classifier $f_{T}$ is designated for the target domain. The architecture of our proposed model is depicted in Fig. 1. Central to our approach is an ensemble model, formulated as:
@@
Here, $w_{i}$ represents the weight parameter for the $i$-th source domain classifier, and $w_{T}$ denotes the weight parameter for the target domain classifier. Crucially, each source domain classifier $f_{S}^{i}$ is pre-trained on its respective domain data, while the target domain classifier $f_{T}$ is dynamically trained and updated with new instances from the target domain as they emerge.
@@
Our ensemble method capitalizes on the strengths of individual classifiers trained on distinct historical data patterns from each source domain, integrating these with the versatile classifier of the target domain. This integration aims to enhance the model’s overall predictive capabilities and accuracy, which is especially beneficial when dealing with limited initial data in the target domain. For weight allocation in our ensemble model, adaptive weights provide a dynamic adjustment mechanism, encompassing two strategies:
Predictive Performance-Adaptive Weighting
This Predictive Performance-Adaptive Weighting (PPAW) strategy involves updating all weights, including those of the target domain classifier and source domain classifiers, solely based on prediction outcomes as shown in Algorithm 1. As new training instances are processed, the model evaluates the accuracy of each classifier, adjusting weights based on their predictive performance: weights of accurate classifiers remain the same while inaccurate classifiers receive a weight penalty, reducing their influence in future predictions.
Input: Pre-trained source models $f_{S}^{1}(\cdot),\ldots,f_{S}^{n}(\cdot)$, Target model $f_{T}(\cdot)$, Target Dataset $X=\{X_{j}\}$ for $j=1,2,\ldots$, Weight penalty $\alpha\in\left(0,1\right)$, Target weight factor $\beta\in\left(0,1\right)$. Initialization: Initialize source model weights $w_{S}^{i}=1.0$ for $i=1,2,\ldots,n$ and target model weight $w_{T}=1.0$. while receive new data point $X_{j}$ do
       if algorithm = PVAW then
@@
       end if
      Normalize weights:
            $w_{S}^{i}\leftarrow\frac{w_{S}^{i}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}},$       $w_{T}\leftarrow\frac{w_{T}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}}$
       Calculate combined probability for final prediction:
@@
             if $f_{S}^{i}(X_{j})$ predicts incorrectly then
                   Adjust weight: $w_{S}^{i}\leftarrow\alpha\times w_{S}^{i}$
@@
      if $f_{T}(X_{j})$ predicts incorrectly then
             Adjust weight: $w_{T}\leftarrow\alpha\times w_{T}$
@@
      Update $f_{T}$ with $X_{j}$;
       Return $prob$ as the predicted probability for $X_{j}$;
@@
Algorithm 1 Predictive Adaptive Weighting Strategies: Performance-Adaptive (PPAW) and Volume-Adaptive (PVAW)
Predictive Volume-Adaptive Weighting
The Predictive Volume-Adaptive Weighting (PVAW) strategy advances beyond the PPAW by innovatively adjusting the weights of classifiers in response to changes in data volume and prediction accuracy. For the target domain classifier, its weight is dynamically modified based on two key factors: the growing volume of target data and its prediction results. As the amount of target domain data increases, the classifier’s weight is proportionally enhanced. This increase is quantified by multiplying the weight with a factor that signifies the quantity of target data, a process facilitated by the variable $j$. The variable $j$ serves as a sequential index for each new data point, effectively tracking the expansion of the target dataset. As more data points are received (i.e., as $j$ increases), the target classifier’s weight is correspondingly adjusted upwards. This adjustment ensures that the weight of the target classifier is attuned to the scale of the available data, enabling the model to progressively prioritize more recent and relevant insights. Simultaneously, the source domain classifiers’ weights are adjusted primarily based on their prediction performance. This dual mechanism—adapting weights for both prediction accuracy and data volume—enables a more dynamic and responsive approach in the PVAW strategy. The target classifier’s weight becomes increasingly significant as more target data becomes available, ensuring that the model’s predictions are continually refined and aligned with the latest trends and patterns in the target domain.
The PPAW and PVAW strategies, detailed in Algorithm 1, aim to optimize the ensemble model’s performance. PPAW modifies response based on classifier accuracy, while PVAW further tailors the target model’s weights to the growing data volume in the target domain. This method ensures the ensemble model stays attuned to the target domain’s evolving dynamics, prioritizing classifiers that resonate with the latest data trends.
We hypothesize that the adaptive nature of learnable weights will demonstrate superior performance compared to static weight assignment strategies. This hypothesis is predicated on the notion that dynamically adjusting weights will allow the ensemble model to evolve in sync with the incoming target data, potentially leading to more accurate and adaptable predictions.

### 07 Apr 2024 v2

@@ Abstract
Transfer learning has become a pivotal technique in machine learning and has proven to be effective in various real-world applications. However, utilizing this technique for classification tasks with sequential data often faces challenges, primarily attributed to the scarcity of class labels. To address this challenge, we introduce Multi-Source Adaptive Weighting (MSAW), an online multi-source transfer learning method. MSAW integrates a dynamic weighting mechanism into an ensemble framework, enabling automatic adjustment of weights based on the relevance and contribution of each source (representing historical knowledge) and target model (learning from newly acquired data). We demonstrate the effectiveness of MSAW by applying it to detect Respiratory Syncytial Virus cases within Emergency Department visits, utilizing multiple years of electronic health records from the University of Pittsburgh Medical Center. Our method demonstrates performance improvements over many baselines, including refining pre-trained models with online learning as well as three static weighting approaches, showing MSAW’s capacity to integrate historical knowledge with progressively accumulated new data. This study indicates the potential of online transfer learning in healthcare, particularly for developing machine learning models that dynamically adapt to evolving situations where new data is incrementally accumulated.
@@
online transfer learning, ensemble method, dynamic weighting mechanism, electronic health record, Respiratory Syncytial Virus case detection
@@ I Introduction
Machine learning has made substantial advancements in recent decades, with its applications spanning a wide range of fields such as computer vision (CV), natural language processing (NLP), speech recognition, and autonomous driving. Despite these achievements, machine learning in biomedicine faces significant challenges, particularly in data collection. The acquisition of labeled data can be very costly or even infeasible due to factors like ethical considerations, patient privacy, and the scarcity of certain diseases. These challenges have led researchers to increasingly rely on utilizing data from related domains that have a more abundant supply of data.
@@
To better leverage knowledge from these related yet distinct domains, transfer learning has emerged as a promising approach to enhance model performance. Fundamental to transfer learning are the concepts of the ‘source domain’ and ‘target domain’. The source domain, which is usually rich in data, serves as the foundation for knowledge extraction. It is characterized by its abundance of well-labeled datasets, which aid models in understanding and encoding complex patterns. In contrast, the target domain, where this pre-learned knowledge is used, often features a more limited dataset. Transfer learning bridges this gap by adapting models trained on the source domain to function effectively in the target domain, thereby helping address the challenges associated with data scarcity.
@@
Traditional transfer learning approaches [1, 2, 3] typically depend on having complete datasets from both the source and target domains at the start of the training process. However, this is often not feasible, especially in dynamic environments where target domain data are collected progressively. To tackle this issue, the concept of online transfer learning [4, 5, 6] has been introduced. This method is designed for scenarios where target domain data are acquired incrementally over time, while source domain data are already accessible before training begins. In such cases, transferring knowledge from the source domain becomes crucial, particularly because the limited initial data in the target domain may be insufficient for effective learning. The extensive and diverse information available from the source domains can significantly compensate for this shortfall, providing an initial knowledge base that the model can build upon as more target domain data becomes available. Therefore, the efficiency and effectiveness of learning in the target domain are greatly enhanced by the transferred knowledge from the source domains.
@@
Online transfer learning entails leveraging knowledge from a static source domain and applying it to an ongoing, evolving target task. The model is incrementally updated as new target data are received. A crucial element of this process is the utilization of a loss function to fine-tune the model in accordance with the latest incoming data, thus ensuring its continued relevance and efficiency. This adaptability renders online transfer learning particularly suitable for real-time applications, effectively handling issues such as concept drift [7] and providing a more versatile approach to knowledge transfer.
@@
In traditional transfer learning, the distinction between various domains is clear and easily identifiable. This clarity allows for straightforward categorization based on common characteristics, such as grouping images by distinguishing features like those of cats versus dogs. However, online transfer learning presents a nuanced challenge as it focuses on the evolution of a single task over time, making the determination of where to apply temporal segmentation less clear. This difficulty arises because the task’s concept remains unchanged, while the data representing the task may significantly change due to temporal factors. This evolution can alter the relevance of past information to current or future tasks, making it difficult to understand how these varying time periods connect and influence one another. Unlike static categorizations, where similarities and differences are often clear and measurable, the dynamic nature of temporal domains means that the relevance of past data to the present task can be fluid and elusive. For instance, in healthcare, temporal segmentation might involve dividing patient data into annual periods to capture seasonal disease trends. With multiple source domains, integrating them optimally into an ensemble model is challenging. A critical question arises: how do we determine which source domains are most relevant to the current period (or target domain) and which are less relevant or unrelated?
@@
Our proposed method employs a sequential data acquisition strategy to tackle this challenge. Consider a healthcare scenario where the task is predicting patient outcomes in a new medical facility. To address initial data scarcity in this facility, our approach significantly relies on pre-trained models derived from source domain data, such as historical patient outcomes from established facilities. These pre-trained source models form the initialization of our ensemble model. This ensemble is adept at autonomously evaluating and determining the relevance of each source model’s learned patterns to the current patient data in the new facility (target model). For instance, a source model trained on a large urban hospital might be initially prioritized for its broad spectrum of cases. As more patient data accumulates in the new facility, the ensemble dynamically adjusts its reliance on this urban model, potentially shifting focus to another source model that aligns more closely with emerging local trends. This strategic utilization of pre-trained source models ensure that the most effective ones are prioritized during the critical early stages of training in the target domain.
@@
As the target domain accumulates more data, the ensemble model modifies its approach. Initially dependent on pre-trained source models, it gradually begins to prioritize the continually updated target model. This change is twofold: a shift in reliance from source to target models and a nuanced adjustment in the weight each model holds in decision-making. This transition enables the model to evolve alongside the growing target domain dataset. The distinct roles of shifting focus and adjusting weights demonstrate the adaptability of our online transfer learning approach, particularly in environments where data are collected over time.
@@
Our framework makes two main contributions. First, we introduce an online multi-source transfer learning architecture for machine learning using data over multiple years. Second, we explore the performance of different ways of generating weights for the ensemble model to improve its accuracy.
@@
We evaluate the effectiveness of our approach using real-world Respiratory Syncytial Virus (RSV) patient encounter data that were collected over multiple years from Emergency Departments (EDs) at University of Pittsburgh Medical Center (UPMC). Our primary target-domain goal is to utilize patients’ signs and symptoms to accurately classify them as either having RSV or not, a critical step towards improving infectious disease case detection using EHRs. RSV is a major cause of respiratory infections that can lead to hospitalizations [8]. The burden of RSV continues to grow, especially in young children aged less than 5 years old[9, 10]. RSV also has been identified as an important disease causing hospitalizations of the elderly and high-risk adults [11]. The disease typically exhibits a seasonal pattern in case numbers [12]. In this study, we segment the RSV case data into distinct domains based on a defined annual cycle, with each domain spanning from June of the previous year to May of the current year. This segmentation strategy is crucial to our framework, aiming to improve RSV case detection accuracy by using data from multiple preceding seasons as source inputs. For instance, we use patient encounter data from June 2011 to May 2018 to develop an ensemble model for classifying encounters from June 2018 to May 2019. The experimental results indicate that our approach improves classification performance, particularly related to public health planning and resource allocation for anticipated RSV seasons.
@@ II Related Work
Traditional transfer learning methodologies typically assume that the source and target domains are predefined, aiming to transfer knowledge from the source to the target domain. Many comprehensive surveys [13, 14, 15, 16, 17] have detailed and categorized a broad range of transfer learning approaches, evaluating factors such as algorithm types and application scenarios. A notable recent development is the growing body of literature reviews focusing on the integration of transfer learning with other advanced machine learning techniques. These include combinations with deep learning [18, 19], reinforcement learning [20, 21], and federated learning [22, 23]. Additionally, some reviews have specifically targeted the application of transfer learning in key areas such as CV [24, 25], NLP [26, 27, 28], healthcare [29, 30, 31, 32, 33], and wireless communication [34, 35], providing insights into the practical implications and utility of transfer learning in these specific fields.
@@
Online transfer learning has emerged to address situations where data are received sequentially and are not available at the start of training. This approach is particularly relevant in dynamic environments where real-time data processing is crucial. Some popular online transfer learning methods [36, 37, 38, 39] have demonstrated excellent performance in many real-world scenarios. Online transfer learning methods can be categorized into data-based methods and model-based methods.
@@
Data-based transfer learning methods involve the transfer of original data to a target domain, allowing the target models to be trained with these augmented data, thereby enriching the available data within the target domain. Online Heterogeneous Knowledge Transition[36] trains the target model using data that co-occur in source and target domains. The co-occurrence data can be a set of text-image pairs, for example, the image includes a red car and the text is ‘red; wheel; transportation’.
@@
Conversely, model-based transfer learning methods focus on firstly building models within the source domain, then adapting both the model structure and parameters for the target domain. Here, ‘model structure’ refers to the architecture or configuration of the model (such as layers in a neural network), while ‘model parameters’ are the specific coefficients that are learned from the source data. The target model leverages insights from the source domain to potentially improve performance. Ensemble models, which combine the outputs or features of multiple models, are commonly employed in model-based methods to further improve accuracy and robustness [37, 38]. Online Heterogeneous Transfer by Hedge Ensemble [39] handles the online heterogeneous transfer problem by updating weights of offline and online classifiers based on the Hedge weighting strategy.
@@
For our online transfer learning algorithm, we used RSV detection as a representative example task. The use of machine learning for RSV detection has been growing. Tso et al. [40] developed an extreme gradient boost decision tree model to detect RSV cases in both emergency department and inpatient settings, using clinical measurements from EHRs, including respiratory rate and peripheral oxygen saturation. In another study, Cord et al. [41] assessed different methodologies such as Logistic Regression, Gaussian Naive Bayes, and Random Forest, ultimately choosing the Light Gradient Boosting Machine algorithm for their RSV detection models.
@@ III-A Problem Formulation
We address the challenge of binary classification in environments where data are dynamic, arriving sequentially and often associated with specific time frames such as epidemiological seasons. Our focus is on creating a method to analyze information collected across seasons, using Bayesian networks as the basis of our classifier training. In practical scenarios, data from each season can be viewed as a separate domain. This temporal aspect presents a distinct challenge, especially at the beginning of a new season when data are scarce. To tackle this problem, we suggest a transfer learning strategy where the current season is the target domain, and data from several past seasons are used as source domains. In transfer learning, research [42] has indicated that simply combining multiple source domains into one is inferior to employing a multi-source strategy. Thus, our method uses data from multiple source domains, represented as $D^{S}=\{D_{S}^{1},D_{S}^{2},\ldots,D_{S}^{n}\}$, where each $D_{S}^{i}$ is data from a specific season, together with a single target domain $D_{T}$. We leverage historical data from past seasons, assuming that patterns and knowledge can effectively inform and improve the classification model for the current season.
@@ III-B Base Model
Our experiment study begins with the implementation of the Naive Bayes classifiers. The effectiveness of Naive Bayes stems from its simplicity, which facilitates tracking model changes across seasons over time, and its computational efficiency, particularly when the features are relatively independent given the class label. An additional benefit of the Naive Bayes model is its statistical efficiency, which allows one to construct a model from a small number of instances, making it particularly well-suited for online learning with only target data.
@@ III-C Online Multi-source Ensemble Model
In our framework, for each source domain $D_{S}^{i}$, we train an offline base model $f_{S}^{i}$ as a classifier on this domain. Similarly, a specific online training classifier $f_{T}$ is designated for the target domain. The architecture of our proposed model is depicted in Fig. 1. Central to our approach is an ensemble model, formulated as:
@@
Here, $w_{i}$ represents the weight parameter for the $i$-th source domain classifier, and $w_{T}$ denotes the weight parameter for the target domain classifier. Each source domain classifier $f_{S}^{i}$ is pre-trained on its respective domain data, while the target domain classifier $f_{T}$ is dynamically trained and updated with new instances from the target domain as they become available.
@@
Our ensemble method capitalizes on the strengths of individual classifiers trained on distinct historical data patterns from each source domain, integrating these with the versatile classifier of the target domain. This integration aims to enhance the model’s overall classification capabilities and accuracy, which is especially beneficial when dealing with limited initial data in the target domain. For weight allocation in our ensemble model, adaptive weights can provide a dynamic adjustment mechanism.
Input: Pre-trained source models $f_{S}^{1}(\cdot),\ldots,f_{S}^{n}(\cdot)$, Target model $f_{T}(\cdot)$, Target Dataset $X=\{X_{j}\}$ for $j=1,2,\ldots$, Weight penalty $\alpha\in\left(1,\infty\right)$, Target weight factor $\beta\in\left(0,1\right)$. Initialization: Initialize source model weights $w_{S}^{i}=1.0$ for $i=1,2,\ldots,n$  target model weight $w_{T}=1.0$, and current data volume in target domain $j$. while receive new data point $X_{j}$ do
       Increase target data volume $j\leftarrow j+1$
@@
       Normalize weights:      $w_{S}^{i}\leftarrow\frac{w_{S}^{i}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}},$ $w_{T}\leftarrow\frac{w_{T}}{\sum_{i=1}^{n}w_{S}^{i}+w_{T}}$
       Get combined probability for final classification:
@@
             if $f_{S}^{i}(X_{j})$ classifies incorrectly then
                   Adjust weight: $w_{S}^{i}\leftarrow w_{S}^{i}\times\frac{\sqrt{j}}{\sqrt{j}+\alpha}$
@@
      if $f_{T}(X_{j})$ classifies incorrectly then
             Adjust weight: $w_{T}\leftarrow w_{T}\times\frac{\sqrt{j}}{\sqrt{j}+\alpha}$
@@
      Update $f_{T}$ with $X_{j}$ Return $prob$ as the classifying probability for $X_{j}$
@@
Algorithm 1 Multi-Source Adaptive Weighting (MSAW)
Our proposed Multi-Source Adaptive Weighting (MSAW) method, detailed in Algorithm 1, involves updating all weights, covering those of the target domain classifier and source domain classifiers. As new training instances are processed, the algorithm assesses the accuracy of each classifier, adjusting weight based on their classifying performance. Weights of accurate classifiers are unchanged while those of inaccurate classifiers are penalized, diminishing their impact on future classifications. MSAW also dynamically adjusts the weights of classifiers based on changes in data volume. As the amount of target domain data increases, the target classifier’s weight is proportionally enhanced. This increase is quantified by multiplying the weight with a factor that signifies the quantity of target data, a process facilitated by the variable $j$. The variable $j$ serves as a sequential index for each new target data point, effectively tracking the expansion of the target dataset. As more target instances are received (i.e., as $j$ increases), the target classifier’s weight is correspondingly adjusted upwards. This adjustment ensures that the weight of the target classifier is attuned to the scale of the available data, enabling the algorithm to progressively prioritize more recent and relevant insights. Simultaneously, the source domain classifiers’ weights are adjusted primarily based on their classification performance. This dual mechanism—adapting weights for both classification accuracy and data volume—enables a more dynamic and responsive approach in the MSAW algorithm. The target classifier’s weight becomes increasingly significant as more target data become available, ensuring that the model’s classifications are continually refined and aligned with the latest trends and patterns in the target domain.
We hypothesize that the adaptive nature of the learnable weight assignment strategy in the proposed algorithm will outperform static weight assignment strategies. This hypothesis rests on the idea that dynamically adjusting weights enables the ensemble model to evolve in sync with incoming target data, potentially leading to more precise and adaptable classifications. The three static weighting assignment strategies for comparison are: (1) The ‘Equal Weights’ method which assigns each source model a constant and equal weight. (2) The ‘Volume-based’ method assigns weights to source models based on the size of their training datasets. (3) The ‘Time-based’ method combines individual models based on their temporal proximity to the testing season (target domain), where ‘time distance’ is defined as the number of seasons between a model’s training data and the testing period.